A stack of $20, $50, and $100 bills were retrieved as part of a police investigation. There were 49 more $50 bills than $100 bills. Also, the number of $20 bills were 3 times the number of $100 bills. If the total value of the money was $6650, find the number of each type of bill.
Answer by mananth(16949)

A stack of $20, $50, and $100 bills were retrieved as part of a police investigation. There were 49 more $50 bills than $100 bills. Also, the number of $20 bills were 3 times the number of $100 bills. If the total value of the money was $6650, find the number of each type of bill.
..
$100 bill = x
$50 bill = x+49
$20 bill = 3x
..
total = $6650
..
100x+50(x+49)+20*3x=6650
...
100x+50x+2450+60x=6650
210x+2450=6650
210x=6650-2450
210x=4200
/210
x=20
....
$100 bill = x = 20
$50 bill = x+49=69
$20 bill = 3x=60
...
m.ananth@hotmail.ca